# Collaborative Document Editing with Multiple Users and AI Agents

### v1

@@ Abstract.
Current AI writing support tools are largely designed for individuals, complicating collaboration when co-writers must leave the shared workspace to use AI and then communicate and reintegrate results. We propose integrating AI agents directly into collaborative writing environments. Our prototype makes AI use transparent and customizable through two new shared objects: agent profiles and tasks. Agent responses appear in the familiar comment feature. In a user study (N=30), 14 teams worked on writing projects during one week. Interaction logs and interviews show that teams incorporated agents into existing norms of authorship, control, and coordination, rather than treating them as team members. Agent profiles were viewed as personal territory, while created agents and outputs became shared resources. We discuss implications for team-based AI interaction, highlighting opportunities and boundaries for treating AI as a shared resource in collaborative work.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Human-centered computing Empirical studies in HCIccs: Computing methodologies Natural language generationccs: Human-centered computing Collaborative and social computing systems and tools
Figure 1. Overview of our prototype: A shared document editor (A) with comments (B), which co-writers can also use to delegate tasks to AI agents (C). Users create custom agents through profiles, via a structured form-like “CV” (D) and/or unstructured notes (E). Agents can be assigned repeatable tasks in a task list (F). Note: The task list (F) opens as a sidebar that replaces the comments sidebar (B), while (D) and (E) are part of a separate full-screen profile view, shown here in parallel for illustration.
[…]
@@ 1. Introduction
Today, co-writers can get AI support, such as with recent writing tools and chatbots based on Large Language Models (LLMs) (Lee et al., 2024). However, these tools are currently designed for individual use. This narrow focus creates friction and overheads for collaboration: Co-writers must leave the shared workspace to interact with AI and later remember, integrate, and communicate the results with their collaborators.
In this paper, we explore a new design direction: direct integration of customisable, shared AI agents into collaborative documents.
Moving beyond the design focus on a single writer’s use of AI is both timely and important: From a usability perspective, AI should support diverse workflows, including the complex dynamics of collaborative writing. Pragmatically, there is growing industry demand to integrate new AI features into widely used tools that already offer real-time collaboration. Finally, through a human-centred lens, if HCI is to shape AI that augments rather than replaces human collaboration, we need design concepts that enhance teamwork rather than push team members out of their shared context.
Integrating AI into collaborative documents presents a challenging design problem: Enabling teams to use AI together demands addressing multiple aspects at the same time, such as collaborative prompting (Han et al., 2024) and task delegation (Tankelevitch et al., 2024), maintaining transparency, and respecting individual and shared needs and preferences (Larsen-Ledet and Borowski, 2021). Prompting AI effectively is already difficult for individuals (Zamfirescu-Pereira et al., 2023) and this challenge is amplified in team contexts, where goals and perspectives must be negotiated (Han et al., 2024). Moreover, no established UI design patterns exist yet for AI in collaborative writing, as reflected in the focus on individual writing found in recent surveys and design theory (Buschek, 2024; Lee et al., 2022; Reza et al., 2025). As recently concluded by Chen et al. (2025), “we have not yet seen AI-assisted writing tools targeting collaborative writing.”
@@
In summary, research has focused on collaborative writing or on writing with AI, but not on the intersection. This motivates our work on these guiding research questions:
@@
  • RQ2: How do co-writers interact with shared AI agents?
To address these questions, we developed a functional prototype of a collaborative text editor that embeds AI agents (Figure 1). Our design introduces two novel shared objects: (1) agent profiles, allowing teams to configure AI personas, and (2) agent tasks, which make AI task delegation explicit and visible to all collaborators. AI responses are integrated via collaborative comments, enabling co-writers to engage with AI output within a familiar UI concept.
We deployed this prototype as a design probe, giving it for a week to 30 participants working in 14 groups. Our analysis of interaction logs and interviews revealed how teams incorporate AI into existing norms of authorship, control, and coordination, rather than treating agents as equal team members. Agent profiles were treated as personal territory, while created agents and their outputs became shared resources. Teams also deliberated over the number of agents. Some preferred the efficiency of a single, general-purpose agent, while others saw multiple role-specific agents as an investment into future prompting and diverse perspectives.
In summary, this paper contributes (1) a UI and interaction design concept and prototype for shared AI in collaborative writing, (2) an empirical study of team-AI interaction behaviour and perception, and (3) design insights for integrating AI into collaborative documents. In a broad view, our findings contribute to the literature on team-AI interaction and writing tools by revealing both opportunities and boundaries for treating AI as a shared resource in collaboration.
@@ 2. Background and Related Work
We discuss related work on collaborative writing with and without AI, writing tools that involve AI, and group interaction with AI.
@@ 2.1.2. Writing support from the crowd and AI
Some research proposed text editors that involve crowd workers (Bernstein et al., 2010; Feldman and McInnis, 2021; Huang et al., 2020, 2025; Kim et al., 2014), where multiple people (and sometimes AI (Huang et al., 2025)) contribute text. However, hese systems still center on one main writer who requests support rather than team collaboration. For instance, in Heteroglossia (Huang et al., 2020) and Inspo (Huang et al., 2025), the writer highlights a span to solicit suggestions from crowd workers or an LLM, which then appear in a sidebar. That is, the interaction design does not involve multiple writers in jointly managing AI input or evaluating its results. Addressing these gaps motivates our design exploration.
@@ 2.3. Collaborative Writing and Adding AI
Next, we reflect on insights into collaborative writing (without AI) and how these motivated the features we explore in this work.
@@ 2.3.1. Bringing AI into the collaborative home of writing
Larsen-Ledet et al. (2020) studied artifact ecologies of collaborative writing, that is, how co-writers use and transition between multiple tools. Based on interviews, they describe the concept of a collaborative home, where writing occurs in a synchronized online platform (e.g. Google Docs, our prototype). This pattern implies that all writers see the same content and tools/UI. In this light, our design explores how we might integrate AI features directly into the collaborative home to reduce transition costs between individual AI tools and the shared document.
@@ 2.4. Interaction with Multiple Agents
Prior work has explored interactive systems with multiple AI agents. Naik et al. (2025) interviewed early adopters to identify use cases, and Lee et al. (2025) introduced a multi-agent system for resolving group conflicts (e.g. competing requests to a care robot). While not focused on writing, both emphasize transparency and coordination. We address these through a shared agent task list and by integrating AI output into collaborative comments.
Klieger et al. (2024) studied collaboration in game development with three predefined agent personas. We extend this by enabling users to create their own agents through a profile UI. Related, Siddiqui et al. (2025) provided six fixed personas for prompting AI (“writer’s friends”), while Benharrak et al. (2024) allowed users to create many such personas. However, both projects did not consider these as agents and thus did not explore variants of initiative or integration with collaborative features, as in our case here.
Finally, we distinguish our work from text generation systems using multiple LLMs (e.g. (Venkatraman et al., 2024)), which exclude human writers. More broadly, in this paper we conceptualize “multi-agent systems” from the user’s perspective – as shared, interactive objects in the UI – rather than as (internal) NLP system architectures.
@@ 2.5. Summary
In summary, prior research remains divided: It has either focused on collaborative writing among multiple people or on writing with AI, but not on the intersection. Existing AI writing tools are overall designed for individual users and do not explore the collaborative dynamics of writing with AI, as also concluded by Chen et al. (2025) recently. To the best of our knowledge, this work is the first to investigate how multiple people work together with multiple shared AI agents within a shared document environment.
@@ 3. Concept and System Characteristics
We describe our overarching design goals and the concrete design features that we implemented in our prototype. For each, we include our motivation based on the literature.
@@ 3.2. Agent Profiles
As one key characteristic, our design explores enabling co-writers to create multiple AI agents as part of their collaborative home (Larsen-Ledet et al., 2020).
@@ 3.2.1. Creating multiple shared agents
Co-writers often take on different roles (Posner and Baecker, 1992; Lowry et al., 2004), motivating support for multiple AI agents with distinct roles. Individuals may also have different needs (Larsen-Ledet and Borowski, 2021); multiple agents let them tailor support accordingly. At the same time, our design keeps the agents visible and editable to everyone to support their use as common objects (Larsen-Ledet et al., 2020).
@@ 3.2.2. Creating agents in structured and unstructured ways
We designed a dedicated UI for agent creation (Figure 2). Profiles can be built through a structured form (CV-like, Figure 2A), an unstructured canvas with notes (Figure 2B), or a mix of both. The structured form is motivated by work on AI personas for writing feedback (Benharrak et al., 2024), which showed that predefined fields (e.g. role, expertise) help users get started. These defaults also reflect findings by Gero et al. (2023) on what writers value as individuality of support actors. The unstructured canvas allows users to add any information they find relevant.
@@ 3.2.3. Suggestions for agent attributes
To help users define agents with relevant details, the structured profile UI offers to generate suggestions (Figure 2C) based on what has already been entered. This design is motivated by findings that prompting can be challenging, in general (Zamfirescu-Pereira et al., 2023), in collaborative settings (Han et al., 2024), and when describing agents (Benharrak et al., 2024).
@@ 3.2.4. Agent presets
We provide presets (Figure 2D): reviewer, idea generator, structure & formatting, and English teacher. Presets help users get started, addressing challenges of prompting (Benharrak et al., 2024; Zamfirescu-Pereira et al., 2023) and metacognitive demands (Tankelevitch et al., 2024). They exemplify roles and tasks that can be delegated to agents. Selecting a preset adds it to the document’s agent list, where it can be edited like any custom agent.
@@ 3.2.5. Generated agent summary as an overview
Our design automatically generates a text summary of each agent (Figure 2E). This is in line with the conclusion in related work (Chen et al., 2025) that AI summaries could help co-writers “sync up” – in our case, keeping an overview and understanding the agents that others have created at a glance.
@@ 3.3. Task List for Agents
Task partitioning, coordination, and delegation are central to collaborative writing (Baecker et al., 1993; Dourish and Bellotti, 1992; Sharples et al., 1993), often supported through materials such as comments and plans (e.g. todo lists) (Bødker et al., 2024). Users in mixed human-AI teams especially value explicit shared goals (Schelble et al., 2022). Moreover, Tankelevitch et al. (2024) identify task delegation as a key metacognitive demand in working with generative AI, calling for designs to support this. Together, these findings motivate our design of an explicit task list for AI agents.
@@ 3.3.1. Creating and editing tasks
Users create tasks in a fold-out sidebar next to the page view (Figure 3A), entering description (instruction/prompt), interaction type, and optionally assigning specific agents. Shortcuts in the top menu (Figure 1, top left) open the task list or add a new task. Existing tasks can be directly triggered, edited, or deleted via buttons (Figure 3C).
@@ 3.3.2. Task initiative
Our design explores two types of task initiative: autonomous and manual. For the autonomous mode, users specify an event to trigger the task (Figure 3B): short intervals, user inactivity, all users offline, document saved, or after several collaborative edits. We informed these options via a survey with academic writers in our network (N=16). In manual mode, tasks run when a user clicks the green “play” button in the list (Figure 3C) or a saved shortcut (see Section 3.3.3 below). Section 4 provides technical details.
@@ 3.3.3. Task shortcuts as tools
Tasks can be saved as shortcuts (Figure 3D), displayed as buttons in a floating toolbar whenever text is selected. This enables repeating frequent tasks without opening the task list. More broadly, making prompts repeatable as UI elements is motivated by recent work such as DirectGPT (Masson et al., 2024), ABScribe (Reza et al., 2024), DynaVis (Vaithilingam et al., 2024), and the toolbar idea by Dang et al. (2022). As shown in Figure 3D, our toolbar also provides four default tools, available even before creating custom shortcuts (Extend, Summarize, Translate, and adding a comment).

### v2

@@ Abstract.
Current AI writing support tools are largely designed for individuals, complicating collaboration when co-writers must leave the shared workspace to use AI and then communicate and reintegrate results. We propose integrating AI agents directly into collaborative writing environments. Our prototype makes AI use visible to all users through two new shared objects: user-defined agent profiles and tasks. Agent responses appear in the familiar comment feature. In a user study (N=30), 14 teams worked on writing projects during one week. Interaction logs and interviews show that teams incorporated agents into existing norms of authorship, control, and coordination, rather than treating them as team members. Agent profiles were viewed as personal territory, while created agents and outputs became shared resources. We discuss implications for team-based AI interaction, highlighting opportunities and boundaries for treating AI as a shared resource in collaborative work.
@@
journalyear: 2026copyright: ccconference: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainbooktitle: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems (CHI ’26), April 13–17, 2026, Barcelona, Spaindoi: 10.1145/3772318.3790648isbn: 979-8-4007-2278-3/2026/04ccs: Human-centered computing Empirical studies in HCIccs: Computing methodologies Natural language generationccs: Human-centered computing Collaborative and social computing systems and tools
Figure 1. Overview of our prototype: A shared document editor (A) with comments (B), which co-writers can also use to delegate tasks to AI agents (C). Users create custom agents through profiles, via a structured form-like “CV” (D) and/or unstructured notes (E). Agents can be assigned to repeatable tasks in a task list (F). Note: This task list (F) opens as a sidebar that replaces the comments sidebar (B), while (D) and (E) are part of a separate full-screen profile view, shown here in parallel for illustration.
@@ 1. Introduction
Today, co-writers can get AI support, such as with recent writing tools and chatbots based on Large Language Models (LLMs) (Lee et al., 2024). However, since these tools are currently designed for individual use, co-writers must leave the shared workspace to interact with AI and later remember, integrate, and communicate the results with their collaborators. Not limited to AI, such tool switches away from collaborative documents are known to be challenging (Larsen-Ledet et al., 2020).
In this paper, we explore a new design direction: direct integration of customisable, shared AI agents into collaborative documents. With “customisable”, we concretely mean allowing users to define their own agents and tasks for those.
Moving beyond the design focus on a single writer’s use of AI is both timely and important: From a usability perspective, AI should support diverse workflows, including the complex dynamics of collaborative writing. Pragmatically, there is growing industry demand to integrate new AI features into widely used tools that already offer real-time collaboration. Finally, through a human-centred lens, if HCI is to shape AI that augments rather than replaces human collaboration, we need designs that enhance teamwork rather than push team members out of their shared context.
Integrating AI into collaborative documents presents a challenging design problem: Enabling teams to use AI together demands addressing multiple aspects at the same time, such as collaborative prompting (Han et al., 2024) and task delegation (Tankelevitch et al., 2024), maintaining transparency, and respecting individual and shared needs and preferences (Larsen-Ledet and Borowski, 2021). Prompting AI effectively is already difficult for individuals (Zamfirescu-Pereira et al., 2023) and this challenge is amplified in team contexts, where goals and perspectives must be negotiated (Han et al., 2024). Moreover, there are currently no established UI design patterns for AI in collaborative writing. This is reflected in recent surveys and design theory that focus on individual writing (Buschek, 2024; Lee et al., 2024; Reza et al., 2025). As recently concluded by Chen et al. (2025), “we have not yet seen AI-assisted writing tools targeting collaborative writing.”
@@
In sum, research has focused on collaborative writing or on writing with AI, but not on the intersection. This motivates our research questions on collaborative use, specifically on how teams experience and work with agents integrated into collaborative writing:
@@
  • RQ2: How do co-writers collaboratively interact with shared AI agents?
To address these questions, we developed a functional prototype of a collaborative text editor that embeds AI agents (Fig. 1). Our design introduces two novel shared objects: (1) agent profiles, allowing teams to configure AI personas, and (2) agent tasks, which make AI task delegation explicit and visible to all collaborators. AI responses are integrated via collaborative comments, enabling co-writers to engage with AI output within a familiar UI concept.
We deployed this prototype as a technology probe (Feng et al., 2025; Hutchinson et al., 2003), giving it for a week to 30 participants working in 14 groups. Our analysis of interaction logs and interviews revealed how teams incorporate AI into existing norms of authorship, control, and coordination, rather than treating agents as equal team members. Agent profiles were treated as personal territory, while created agents and their outputs became shared resources. Teams also deliberated over the number of agents. Some preferred a single, general-purpose agent because they perceived it as efficient, while others saw multiple role-specific agents as an investment into future prompting and diverse perspectives.
In summary, this paper contributes (1) a UI and interaction design concept and prototype for shared AI agents in collaborative writing, (2) an empirical study of team-AI interaction behaviour and perception, and (3) design insights for integrating AI into collaborative documents. In a broad view, our findings contribute to the literature on team-AI interaction and writing tools by revealing both opportunities and boundaries for treating AI as a shared resource in collaboration.
@@ 2. Background and Related Work
We discuss related work on collaborative writing, writing tools that involve AI, and group interaction with AI.
@@ 2.1.2. Writing support from the crowd and AI
Some research proposed text editors that involve crowd workers (Bernstein et al., 2010; Feldman and McInnis, 2021; Huang et al., 2020, 2025; Kim et al., 2014), where multiple people (and sometimes AI (Huang et al., 2025)) contribute text. However, these systems still focus on one main writer who requests support rather than team collaboration. For instance, in Heteroglossia (Huang et al., 2020) and Inspo (Huang et al., 2025), the writer highlights a span to request suggestions from crowd workers or an LLM, which then appear in a sidebar. That is, the interaction design does not involve multiple writers in jointly managing AI input or evaluating its results. Addressing these gaps motivates our design exploration.
@@ 2.3. Collaborative Writing and Adding AI
Next, we reflect on insights into collaborative writing (without AI) and how these motivated the features we explore.
@@ 2.3.1. Bringing AI into the collaborative home of writing
The notion of artifact ecologies deals with how collaborative writers involve multiple tools and transition between them: Larsen-Ledet et al. (2020) found that such transitions introduce challenges, such as “When writers work outside the common documents and applications, their co-writers lose sense of the progress and who is working on what.” This maps well to the situation with AI tools today, which are separated from shared documents. Larsen-Ledet et al. (2020) also describe the concept of a collaborative home, where writing occurs in a synchronized platform such that all writers see the same content and tools/UI. Motivated by these prior observations on transition costs and shared tools, our design explores how teams collaborate and use AI integrated directly into the collaborative home.
@@ 2.3.3. Supporting idiosyncratic preferences and common objects, also for AI
Table 1. Our main design elements, compared to related work. The overarching novelty is that we allow for collaboratively creating and interacting with agents and tasks, directly within a shared document editor, including in comments.
Four-column table summarising design elements, their use in related work, our design decisions, and motivations, also in the context of a technology probe study.
Design element Examples in related work and products In our design Motivation for our design Motivation in context of our technology probe study
Multiple agent personas in the context of writing Benharrak et al. (2024): A single user can define multiple agents with a structured UI and request feedback on text, shown in a sidebar. Siddiqui et al. (2025): A single user can request feedback on text from pre-defined agents, shown in a pop-up. Multiple users can define multiple agents, separately or collaboratively, by editing agents created by others, both via structured (“CV”) an unstructured (“notes”) UIs. Support collaborative creation of shared, custom agents for collaborative use. Give participants the opportunity to experience and explore creating and sharing agents in a collaborative document context.
Repeatable task list Broadly related designs include apps for end-user development of automation rules (e.g. IFTTT (IFTTT, 2025)) and “prompt libraries” such as in Claude (19), which list prompts (text snippets) for users to copy into a chatbot conversation. Users create tasks (via free text prompts) available for editing and use by all collaborators. They can assign tasks to agents, regardless of who had created the agent. Tasks are persisted and repeatable via automated triggers or clicking a button. Support collaborative creation and use of tasks for agents. Support both manual and automated task triggers. Give participants the opportunity to experience and explore creation and shared use of agent tasks in a collaborative document context, and allow them to experiment with varying degrees of agent initiative.
Comments Established feature in widely used products (e.g. Microsoft Word, Google Docs). Users can write comments that tag any agent(s) to trigger them to respond. User can also ask one agent to respond to another one. Agents can also add new comments on their own in response to document tasks. Support collaborative interaction with shared agents by integrating it into a UI element already used by teams for (human-to-human) collaboration. Give participants the opportunity to experience and explore interaction with agents (with varying degrees of initiative) in the collaborative writing context.
@@ 2.4. Interaction with Multiple Agents
Prior work has explored interactive systems with multiple AI agents. Naik et al. (2025) interviewed early adopters to identify use cases, and Lee et al. (2025) introduced a multi-agent system for resolving group conflicts. While not focused on writing, both emphasize transparency and coordination. We address these through a shared agent task list and by integrating AI output into collaborative comments.
Klieger et al. (2024) studied collaboration in game development with three predefined agent personas. We extend this by enabling users to create their own agents through a profile UI. Related, Siddiqui et al. (2025) provided six fixed personas for prompting AI (“writer’s friends”), while Benharrak et al. (2024) allowed users to create many such personas. However, both projects did not consider these as agents and thus did not explore variants of initiative or integration with collaborative features.
Finally, we distinguish our work from text generation systems using multiple LLMs (e.g. (Venkatraman et al., 2024)), which exclude human writers. We conceptualize “multi-agent systems” from the user’s perspective – as shared, interactive objects in the UI.
@@ 2.5. Summary
Prior research has either focused on collaborative writing among multiple people or on writing with AI, but not on the intersection. Existing AI writing tools are overall designed for individual users and do not explore the collaborative dynamics of writing with AI, as also concluded by Chen et al. (2025). To the best of our knowledge, this work is the first to investigate how multiple people work together with multiple shared AI agents within a shared document environment. Table 1 summarises our key design elements in comparison to related appearances. The next section describes our design in detail.
@@ 3. Concept and System Characteristics
We describe our overarching design goals and motivate the features implemented in our prototype.
@@ 3.2. Agent Profiles
Our design explores enabling co-writers to create multiple AI agents as part of their collaborative home (Larsen-Ledet et al., 2020).
@@ 3.2.1. Creating multiple shared agents
Co-writers often take on different roles (Posner and Baecker, 1992; Lowry et al., 2004), motivating support for multiple agents with distinct roles. Multiple agents also let users tailor support according to different needs (Larsen-Ledet and Borowski, 2021). Our design keeps agents visible and editable to everyone to support their use as common objects (Larsen-Ledet et al., 2020).
@@ 3.2.2. Creating agents in structured and unstructured ways
In our UI for agent creation (Fig. 2), profiles can be defined in structured form (CV-like, Fig. 2A), on an unstructured canvas with notes (Fig. 2B), or by mixing both. The structured form is motivated by work on AI personas (Benharrak et al., 2024), which showed that predefined fields help users get started. The unstructured canvas allows users to add any information.
@@ 3.2.3. Suggestions for agent attributes
To help users define agents, the structured profile UI offers suggestions (Fig. 2C), motivated by findings that prompting can be challenging, in general (Zamfirescu-Pereira et al., 2023), in collaborative settings (Han et al., 2024), and when describing agents (Benharrak et al., 2024).
@@ 3.2.4. Agent presets
We provide presets (Fig. 2D): reviewer, idea generator, structure & formatting specialist, and English teacher. Presets help users get started, addressing challenges of prompting (Benharrak et al., 2024; Zamfirescu-Pereira et al., 2023) and metacognitive demands (Tankelevitch et al., 2024). Selecting a preset adds it to the document’s agent list, where it can be edited.
@@ 3.2.5. Generated agent summary as an overview
Our design automatically generates a summary of each agent (Fig. 2E), in line with the conclusion in related work (Chen et al., 2025) that AI summaries could help co-writers “sync up” – in our case understanding agents created by others.
@@ 3.3. Task List for Agents
Task partitioning, coordination, and delegation are central to collaborative writing (Baecker et al., 1993; Dourish and Bellotti, 1992; Sharples et al., 1993), supported through materials such as comments and plans (e.g. todo lists) (Bødker et al., 2024). Users in mixed human-AI teams especially value explicit shared goals (Schelble et al., 2022). Tankelevitch et al. (2024) identify task delegation as a metacognitive demand in working with generative AI, calling for designs to support this. These findings motivate our explicit task list for agents.
@@ 3.3.1. Creating and editing tasks
Users create tasks in a fold-out sidebar next to the page (Fig. 3A), entering description (prompt), interaction type, and optionally assigning specific agents. Shortcuts in the top menu (Fig. 1, top left) open the task list or add a new task. Existing tasks can be directly triggered, edited, or deleted via buttons (Fig. 3C).
@@ 3.3.2. Task initiative
Our design explores two types of task initiative: autonomous and manual. For the autonomous mode, users specify an event to trigger the task (Fig. 3B): short intervals, user inactivity, all users offline, document saved, or after several collaborative edits. We informed these options via a survey with academic writers (N=16, see Appendix A). In manual mode, tasks run when a user clicks the green “play” button in the list (Fig. 3C) or a saved shortcut (see Section 3.3.3).
@@ 3.3.3. Task shortcuts as tools
Tasks can be saved as shortcut buttons (Fig. 3D) in a floating toolbar, shown whenever text is selected. Thus, frequent tasks can be repeated without opening the list. Making prompts repeatable as UI elements is motivated by DirectGPT (Masson et al., 2024), ABScribe (Reza et al., 2024), DynaVis (Vaithilingam et al., 2024), and the toolbar idea by Dang et al. (2022). As shown in Fig. 3D, our toolbar also provides four default tools (Extend, Summarize, Translate, and adding a comment).